A Bibliography of Books about the Internet and Networking: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

05 May 2025 Version 1.06

Title word cross-reference

Administration [FM20]. Bitcoin [vO21]. Computer [vO21]. Course [MFD20]. Creation [FM20]. Decisions [SB24]. Digital [SB24]. Finnish [Ano20]. First [MFD20]. Forks [SB24]. Freedom [Fed24]. History [Ano20, SB24]. Identifiers [FM20]. Internet [Ano20, Fed24, SB24, vO21]. jewels [vO21]. Key [SB24]. malware [vO21]. Network [MFD20]. Open [Fed24]. Restoring [Fed24]. Road [SB24]. Safeguarding [Fed24]. Science [MFD20]. Securing [Fed24]. security [vO21]. tools [vO21]. Unique [FM20].

References

Anonymous:2020:HFI

[Ano20] Anonymous. History of the Finnish Internet. Web site, November 30, 2020. URL https://

siy.fi/history-of-the-finnish-
internet/.

FCC:2024:SSO

[Fed24] Federal Communications Commission. Safeguarding and securing the Open Internet and restoring Internet freedom. Ruling FCC 24-52, Federal Communications Commission, Washington, DC 20554, USA, May 7, 2024. 513 (final 82+ missing: should

be 595+) pp. URL https://docs.
fcc.gov/public/attachments/FCC-
24-52A1.pdf.

Fidler:2020:CAU

[FM20] Bradley Fidler and Russ Mundy.

The creation and administration of unique identifiers, 1967–2017. Internet Corporation for Assigned Names and Numbers (ICANN) report, ????, November 2020. 118

pp. URL https://www.icann.org/
en/system/files/files/creation-
administration-unique-identifiers-
1967-2017-18nov20-en.pdf.

Menczer:2020:FCN

[MFD20] Filippo Menczer, Santo Fortunato, and Clayton A. Davis. A First Course in Network Science. Cambridge University Press, Cambridge, UK, 2020. ISBN 1-108-47113-7 (hardcover), 1-108-65394-4 (e-book). xv + 258 pp. LCCN TK5105.5 .M46 2020, T57.85 .M46 2020.

Shackelford:2024:FDR

[SB24] Scott J. Shackelford and Scott O. Bradner. Forks in the Digital Road: Key Decisions in the History of the Internet. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, April 2024. ISBN 0-19-761776-X (hardcover), 0-19-761777-8 (paperback), 0-19-761778-6 (e-book), 0-19-761779-4 (e-book), 0-19-761780-8 (ebook). xxiv + 207 pp. LCCN TK5105.875.I57 .S533 2024. Foreword by Vinton G. Cerf.

vanOorschot:2021:CSI

[vO21] Paul C. van Oorschot. Computer security and the Internet: tools and jewels from malware to Bitcoin. Information security and cryptography. Springer, Cham, Switzerland, second edition, 2021. ISBN 3-030-83410-7 (hardcover), 3-030-83411-5 (e-book). xxix + 446 pp. LCCN QA76.9.A25 V66 2021.